how many kgs of tea at ₹13per kg must be mixed with 100 kg of tea at ₹17.50 per kg,so that a profit of 40% may be made by selling the mixture at ₹21 per kg?
Answer by Boreal(15235)

First, find the cost price or x
40% of x +x=21
1.4x=21
x=15 that is the cost to produce
y kg of tea at 13/kg
100 kg of tea at 17.50/kg
100+y kg of tea at 15/kg, cost price
so 13y+1750=(100+y)15=1500+15y
-2y=-250
y=125 kg of tea at 13/kg or 1625
100kg of tea at 17.50 or 1750
This is 225 kg costing 3375 or 15/kg